Grade 10Algebra

If x3-x2-ax+b is divisible by x2-x.
Writethe values of a and b.

karthikeya
10 Years agoGrade 10

1 Answer

Vikas TU
10 Years ago
If x3-x2-ax+b is divisible by x2-x. then the remainder will be zero and the roots of x^2-x will satify the cubic polynomial therfore,
From Remainder Theorem,
Roots for x^2 – x = 0
are:
x(x -1) = 0
x = 0  or  x = 1
 
f(x) =  x3-x2-ax+b
f(0) = b = 0
and 
f(1) = 0 = >  a = b= 0